OpenStudy (marcelie):

help please number 6

OpenStudy (marcelie):

the curve y=x^2 , 0</ x </ 1 is rotated about the y axis. find the area of the resulting surface

OpenStudy (agent0smith):

You should be able to do this... start with a diagram.

OpenStudy (marcelie):

this is where i left of .

OpenStudy (danjs):

yeah looks alright so far.. you want to integrate over x=0 to x=1 y = x^2 y ' = 2x dx \[\large 2 \pi \int\limits_{0}^{1}x \sqrt{1+4x^2}dx\]

OpenStudy (agent0smith):

Sub u = 1+ 4x^2 du = 8x dx

OpenStudy (danjs):

yes, also make sure to change the bounds, x from 0 to 1 u = 1 + 4x^2 becomes u from 1 to 5

OpenStudy (marcelie):

so then i got this \[\frac{ \pi }{ 6 }(1+4x^2)^3/2 ...from 0\to1\]

OpenStudy (danjs):

\[\large 2\pi \int\limits_{0}^{1}x \sqrt{1+4x^2}dx\] if u = 1+4x^2 du = 8x dx so x*dx = 1/8 du \[\large 2\pi*\int\limits_{1}^{5}\sqrt{u}*\frac{ 1 }{ 8 }du\]

OpenStudy (danjs):

notice also it is from u=1 to u=5 now

OpenStudy (marcelie):

oh riightt

OpenStudy (danjs):

\[= \large \frac{ 2\pi }{ 8 }*\frac{ 2 }{ 3 }*u^{\frac{ 3 }{ 2} }\] from u=1 to u=5